What is the second harmonic for an organ pipe with one open end whose length is 1.13 m and is being played on a day when the temperature is 25.891 oC?
Answer #1

speed of Sound V = 331 +0.61T Vs = 331 + 0.61 425.891) now, V = 346.79 m/s and harmonic freq. L=1 > d=26= 2x1.13 de 2.26m i f